OpenStudy (help!!!!):

determine the value of a so that the graph of the equation passes through the given point

OpenStudy (help!!!!):

3x+ay=15 (-2,3)

OpenStudy (anonymous):

That's an equation of a line. So rearrange the equation ay=15-3x y=15/a -(3x)/a y=(15-3x)/a Substitute the point (-2,3) 3=(15-(3)(-2))/a simplify 3=21/a Solve for a 21/3=a 7=a